OpenStudy (anonymous):

Need help now please! Inequalities: Maria solved the equation -5 1/8n = -9 3/4. Her answer was n = 1 37/41. She is fairly confident about her answer, but wants to do a quick check of her solution using estimation. Which of the following shows a good check of her answer? -6(2) = -12 -5(-2) = -10 -5(2) = -9 -5(2) = -10 thank u!! :)

OpenStudy (ranga):

\[\Large -5\frac{ 1 }{ 8 }n = -9\frac{ 3 }{ 4 } \\ n = 1\frac{ 37 }{ 41 }\] Are those the numbers?

OpenStudy (anonymous):

Sorry,the page froze up on me! Yes, those are the numbers :)

OpenStudy (ranga):

-5 and 1/8 can be rounded down to -5 -9 and 3/4 can be rounded up to -10 n = 1 and 37/41 can be rounded up to n = 2 Therefore, (-5)(2) = -10

OpenStudy (anonymous):

thanks a million :)

OpenStudy (ranga):

you are welcome.